# An extension of Steinberg’s Theorem to biquotient pairs of subgroups

### v2

@@ 1.3 Connected groups
We now restrict our attention to connected compact $G$. Recall that all maximal tori of $G$ are conjugate, and that any element of $G$ is contained in a maximal torus [Adams:LieLectures, Corollaries 4.22 & 4.23]. Given a maximal torus $T\subseteq G$, the Weyl group $\mathrm{W}:=\mathrm{W}(G,T)$ is defined as $\mathrm{W}:=\mathrm{N}_{G}T/T$. This is always a finite group. It acts on $\mathrm{R}T$ by conjugation, and we denote by $(\mathrm{R}T)^{\mathrm{W}}$ the subring of fixed points under this action.
@@
For the identification of $\mathrm{R}G$ with $(\mathrm{R}T)^{\mathrm{W}}$, see for example [Adams:LieLectures, Theorem 6.20]. For the finiteness of $i^{*}$, see Lemma 7 above. Alternatively, use the identification of $\mathrm{R}G$ with $(\mathrm{R}T)^{\mathrm{W}}$ and [GoertzWedhorn:AGI]*Proposition 12.27 (4).
@@
Let us write $\mathfrak{I}G$ for the kernel of the rank homomorphism $\operatorname{rank}\colon\mathrm{R}G\to\mathbb{Z}$. Clearly, $\mathfrak{I}T$ is a prime ideal of $\mathrm{R}T$ lying over $\mathfrak{I}G$, and it is $\mathrm{W}$-invariant. So Lemma 8 shows that $(\mathrm{R}T)_{\mathfrak{I}G}$ is a local ring with maximal ideal $\mathfrak{I}T\cdot(\mathrm{R}T)_{\mathfrak{I}G}$.
@@
For any compact connected Lie group $G$ with maximal torus $i\colon T\hookrightarrow G$, we have a morphism of $RG$-modules $i_{*}\colon RT\to RG$ such that $i_{*}i^{*}=\operatorname{id}$.
@@
See [Atiyah:Bott]*Rmk 1 after Prop 4.9 or [EastwoodSalomon, § 3], in particular Lemma 3.4. ∎
@@
$\dim\mathrm{R}G=\dim(\mathrm{R}G/\mathfrak{p})+\dim(\mathrm{R}G)_{\mathfrak{p}}.$
@@
    $R$ is catenary, i. e. all saturated chains of prime ideals between two given prime ideals in $R$ have equal length, and
@@
    $R$ is equicodimensional, i. e. all its maximal ideals have equal height.
We now verify these two properties for $R=\mathrm{R}G$. Property (i) clearly holds as $\mathrm{R}G$ is an integral domain (Lemma 8 (a)). Property (ii) holds for $\mathrm{R}G$ as $\mathbb{Z}$ is universally catenary, as is any Dedekind domain, and as $\mathrm{R}G$ is finitely generated over $\mathbb{Z}$ (Lemma 6 (b)). Property (iii) holds for $R=\mathrm{R}G$ in view of the following general fact: any integral domain $R$ finitely generated over a Noetherian Jacobson universally catenary and equicodimensional domain $S$ is again equicodimensional [EGA4, Proposition 10.6.1 (i)]. This applies in particular to $S=\mathbb{Z}$, and more generally to any Dedekind domain $S$ with infinitely many maximal ideals. ∎
We include here another lemma that we will use in conjunction with Lemma 11.
@@
$\dim S_{\mathfrak{p}}-\dim(S/\mathfrak{b})_{\mathfrak{p}}=\dim R_{\mathfrak{p}}-\dim(R/\mathfrak{p}^{\prime})_{\mathfrak{p}}$
@@
For the ideal $\mathfrak{a}:=R\cap\mathfrak{b}$, we have a finite extension $R/\mathfrak{a}\hookrightarrow S/\mathfrak{b}$, and hence a finite extension $(R/\mathfrak{a})_{\mathfrak{p}}\hookrightarrow(S/\mathfrak{)}_{\mathfrak{p}}$ for any prime ideal $\mathfrak{p}\subset R$. In particular, $(\dim R/\mathfrak{a})_{\mathfrak{p}}=(\dim S/\mathfrak{b})_{\mathfrak{p}}$, and similarly $\dim S_{\mathfrak{p}}=\dim R_{\mathfrak{p}}$. Next, note that $\mathrm{Supp}_{R}(S/\mathfrak{b})=\mathrm{Supp}_{R}(R/\mathfrak{a})=\{\mathfrak{p}\in\operatorname{Spec}(R)\mid\mathfrak{a}\subseteq\mathfrak{p}\}.$ Indeed, the first equality again follows from the observation that $(R/\mathfrak{a})_{\mathfrak{p}}\hookrightarrow(S/\mathfrak{b})_{\mathfrak{p}}$ is an extension, and the second equality is elementary. So the given prime ideal $\mathfrak{p}\in\mathrm{Supp}_{R}(S/\mathfrak{b})$ contains $\mathfrak{a}$, and $\dim(R/\mathfrak{a})_{\mathfrak{p}}=\dim(R/\mathfrak{p}^{\prime})_{\mathfrak{p}}$ for some prime ideal $\mathfrak{p}^{\prime}$ of $R$ minimal among the primes satisfying $\mathfrak{a}\subseteq\mathfrak{p}^{\prime}\subseteq\mathfrak{p}$. In particular, $\mathfrak{p}^{\prime}\in\mathrm{Supp}_{R}(S/\mathfrak{b})$ and $\mathfrak{p}^{\prime}\subseteq\mathfrak{p}$. ∎
@@ 1.4 Good groups and Steinberg’s Theorem
    The representation ring $\mathrm{R}G$ is a tensor product of a polynomial ring and a Laurent ring, i. e. isomorphic to the ring $\mathbb{Z}[x_{1},\dots,x_{m},y_{1}^{\pm 1},\dots,y_{n}^{\pm 1}]$ for certain $m$ and $n$.
@@
To avoid endlessly repeating these somewhat intricate conditions, we simply define a good Lie group to be a compact connected Lie group $G$ that satisfies the equivalent conditions of Theorem 13. The result stated as Steinberg’s Theorem is just the implication $(a\Rightarrow c)$ of this theorem combined with the following observation:
@@
Suppose $G$ is compact connected with torsion-free fundamental group. Writing $G$ as $(G^{\prime}\times T)/\Gamma$ as above, we find that the fundamental group of $G^{\prime}$ is also torsion-free. As $G^{\prime}$ is semisimple, this implies that $G^{\prime}$ is simply-connected, and hence a direct product of simply-connected simple groups. So $G$ satisfies condition (a) of Theorem 13. ∎
Characterization (b) in Theorem 13 implies:
@@
For good $G$, $\mathrm{R}G$ is a localization of a polyonimal ring over the integers, which is clearly noetherian and of finite global dimension. ∎
Characterization (b) also implies that, for good $G$, the notions of “projective dimension” and “weak/flat dimenension” are equivalent for finitely generated $\mathrm{R}G$-modules. Indeed:
@@
The implication $(a\Rightarrow b)$ of Theorem 13 is fairly elementary, and the implication $(b\Rightarrow c)$, which is the main ingredient in the result stated as Steinberg’s Theorem in the introduction, was in fact proved slightly earlier by Pittie [Pittie:Homogeneous], modulo the above generalization of the Quillen–Suslin theorem. Thus, from today’s vantage point, the main contribution of Steinberg is a proof of the opposite implication: a precise characterization of those compact Lie groups that have the desired property (c). Moreover, Steinberg’s proof includes an explicit basis of $\mathrm{R}H$ as an $\mathrm{R}G$-module.
We can easily sketch Pittie’s short argument for the implication $(b\Rightarrow c)$ based on the results and observations already included above: For a maximal torus $T$ in a good group $G$, both $\mathrm{R}G$ and $\mathrm{R}T$ are regular noetherian rings (Lemma 15). As $\mathrm{R}T$ is in any case finitely generated as an $\mathrm{R}G$-module (Lemma 7), it thus follows from [Serre:LA, Chapter III.D, § 5, Theorem 13 (b)] that $\mathrm{R}T$ is a projective $\mathrm{R}G$-module. For a general subgroup of maximal rank $H$, we may assume that $H$ contains $T$, and use holomorphic induction (Theorem 10) to deduce that $\mathrm{R}H$ is a direct summand of $\mathrm{R}T$. In particular, $\mathrm{R}H$ is also projective as an $\mathrm{R}G$-module. Finally, as we have just seen in Lemma 16, this is today known to imply that $\mathrm{R}H$ is free.
While Theorem 13 only deals with subgroups of maximal rank, the following consequence for subgroups of arbitrary ranks was proved in [SCSQ, Corollary 3.10]:
@@
We do not reproduce the short proof of Corollary 18 here, but we note that besides holomorphic induction (Theorem 10) the main ingredient is a change-of-rings theorem for Tor. We record here two such theorems that we will have occassion to use later.
@@
This is a special case of [CartanEilenberg, Chapter IX, Theorem 2.8]. Cartan and Eilenberg work with non-commutative $K$-projective algebras $\Lambda$, $\Gamma$, $\Sigma$ over a commutative ring $K$. We have specialized the statement to the case of $\mathbb{Z}$-free commutative $\mathbb{Z}$-algebras. Note that the isomorphism $r$ of [CartanEilenberg, Proposition 2.1], on which the theorem relies, is an isomorphism of $\Sigma\otimes\Gamma$-algebras in this commutative situation. ∎
When $\Lambda$ and $\Gamma$ are free as abelian groups, Theorem 19 can be seen as a special case of Theorem 20.
@@ 1.5 Supports
In [Segal:RG], Segal analyses the prime spectrum $\operatorname{Spec}(\mathrm{R}G)$ of an arbitrary compact Lie group $G$. In particular, Segal shows that, for each prime ideal $\mathfrak{p}\subset\mathrm{R}G$, there is a subgroup $S_{\mathfrak{p}}$, unique up to conjugation, which is minimal with respect to the property that $\mathfrak{p}$ is the preimage of a prime ideal of $\mathrm{R}(S_{\mathfrak{p}})$ under the restriction $\mathrm{R}G\to\mathrm{R}S_{\mathfrak{p}}$. This subgroup $S_{\mathfrak{p}}$, or rather any subgroup in this conjugacy class, is referred to as the support of $\mathfrak{p}$. The following proposition explains, in particular, how Segal’s “supporting subgroups” and the usual notion of “prime ideals supporting a module” are related.
@@
    $\dim(\mathrm{R}G/\mathfrak{p})\leq\operatorname{rank}H+1.$
@@
Claim (a) is clear from the equivalence $(c\Leftrightarrow d)$ of Proposition 21.
The first equality in Claim (b) holds in general for finitely generated modules over a commutative ring. See [AtiyahMacDonald, Chapter 3, Exercise 19 (iv)] for the general fact, and Lemma 7 for finite generation in the case at hand. The second equality in Claim (b) again follows from the equivalence $(c\Leftrightarrow d)$ of Proposition 21.
For claim (c), recall from [Segal:RG] that the support $S_{\mathfrak{p}}$ is “cyclic”, i. e. containes an element $s$ whose powers are dense in $S_{\mathfrak{p}}$. This element $s$ is contained in some maximal torus of $G$, so we may as well assume $s\in T$ and hence $S_{\mathfrak{p}}\subseteq T$. The ideal $\mathfrak{p}\subset\mathrm{R}G$ is thus the inverse image of an ideal $\mathfrak{p}_{0}\subset\mathrm{R}(S_{\mathfrak{p}})$ under a restriction map that factors as
@@
Recall that any inclusion of a subgroup into a torus induces a surjection on representation rings (Lemma 4 (a)). Writing $\mathfrak{p}_{T}\subset\mathrm{R}T$ for the intermediate preimage of $\mathfrak{p}_{0}$, we obtain induced morphisms of quotient rings as follows:
@@
Here, the first map is a finite extension (see Lemma 8), so $\dim\mathrm{R}G/\mathfrak{p}=\dim\mathrm{R}T/\mathfrak{p}_{T}$, and we deduce $\dim(\mathrm{R}G/\mathfrak{p})\leq\dim(\mathrm{R}S_{\mathfrak{p}})$. Moreover, by Lemma 6 (c), $\dim(\mathrm{R}S_{\mathfrak{p}})=\operatorname{rank}(S_{\mathfrak{p}})+1$, and as $S_{\mathfrak{p}}$ is conjugate to a subgroup of $H$ we have $\operatorname{rank}(S_{\mathfrak{p}})\leq\operatorname{rank}H$.
@@
Let us again write $\mathfrak{p}_{T}$ for the intermediate preimage on $\mathrm{R}T$, which by construction lies over $\mathfrak{p}$. The restriction $RT\twoheadrightarrow\mathrm{R}(\mathrm{Z}G)$ clearly commutes with the action of the Weyl group on $\mathrm{R}T$, so the Weyl group acts trivially on $\mathfrak{p}\in\operatorname{Spec}(\mathrm{R}T)$. It therefore follows from Lemma 8 (b) that $\mathfrak{p}_{T}$ is the only ideal over $\mathfrak{p}$. ∎
Example 23.
Recall the defition of the prime ideals $\mathfrak{I}G$ and $\mathfrak{I}_{p}G$ from Example 9. If $(G,H_{1},H_{2})$ satisfy the strict biquotient condition, then the intersection $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$ consists exclusively of the primes $\mathfrak{I}G$ and $\mathfrak{I}_{p}G$ for the various primes $p\in\mathbb{Z}$. Indeed, proposition (b) identifies this intersection with $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(1)$, and by the equivalence $(b\Leftrightarrow c)$ of Proposition 21 this latter support consists precisely of the primes containing $\mathfrak{I}G$.

### v3

@@ 1.3 Connected groups
We now restrict our attention to compact connected $G$. Recall that all maximal tori of $G$ are conjugate, and that any element of $G$ is contained in a maximal torus [Adams:LieLectures, Corollaries 4.22 & 4.23]. Given a maximal torus $T\subseteq G$, the Weyl group $\mathrm{W}=\mathrm{W}(G,T)$ is defined as $\mathrm{W}:=(\mathrm{N}_{G}T)/T$. This is always a finite group. It acts on $\mathrm{R}T$ by conjugation, and we denote by $(\mathrm{R}T)^{\mathrm{W}}$ the subring of fixed points under this action.
@@
For the identification of $\mathrm{R}G$ with $(\mathrm{R}T)^{\mathrm{W}}$, see for example [Adams:LieLectures, Theorem 6.20]. For the finiteness of $i^{*}$, see Lemma˜7 above. Alternatively, use the identification of $\mathrm{R}G$ with $(\mathrm{R}T)^{\mathrm{W}}$ and [GoertzWedhorn:AGI]*Proposition 12.27 (4).
@@
Let us write $\mathfrak{I}G$ for the augmentation ideal, i. e., for the kernel of the rank homomorphism $\operatorname{rank}\colon\mathrm{R}G\to\mathbb{Z}$. Clearly, $\mathfrak{I}T$ is a prime ideal of $\mathrm{R}T$ lying over $\mathfrak{I}G$, and it is $\mathrm{W}$-invariant. So Lemma˜9 shows that $(\mathrm{R}T)_{\mathfrak{I}G}$ is a local ring with maximal ideal $\mathfrak{I}T\cdot(\mathrm{R}T)_{\mathfrak{I}G}$.
@@
For any compact connected Lie group $G$ with maximal torus $i\colon T\hookrightarrow G$, we have a morphism of $\mathrm{R}G$-modules $i_{*}\colon\mathrm{R}T\to\mathrm{R}G$ such that $i_{*}i^{*}=\operatorname{id}$.
@@
See [Atiyah:Bott]*Remark 1 after Proposition 4.9 or [Segal:RG, § 4], where $i_{*}$ is written as $i_{c}$ and the identity $i_{c}i^{*}=\operatorname{id}$ follows from Proposition 4.1 (iv). ∎
@@
$\dim\mathrm{R}G=\dim{\mathrm{R}G/\mathfrak{p}}+\dim{(\mathrm{R}G)_{\mathfrak{p}}}.$
@@
    $R$ is catenary, i. e., all saturated chains of prime ideals between two given prime ideals in $R$ have equal length, and
@@
    $R$ is equicodimensional, i. e., all its maximal ideals have equal height.
We now verify these two properties for $R=\mathrm{R}G$. Property (i) clearly holds as $\mathrm{R}G$ is an integral domain (Lemma˜9 (a)). Property (ii) holds for $\mathrm{R}G$ as $\mathbb{Z}$ is universally catenary, as is any Dedekind domain, and as $\mathrm{R}G$ is finitely generated over $\mathbb{Z}$ (Lemma˜6 (b)). Property (iii) holds for $R=\mathrm{R}G$ in view of the following general fact: any integral domain $R$ finitely generated over a Noetherian Jacobson universally catenary and equicodimensional domain $S$ is again equicodimensional [EGA4, Proposition 10.6.1 (i)]. This applies in particular to $S=\mathbb{Z}$, and more generally to any Dedekind domain $S$ with infinitely many maximal ideals. ∎
We include here another lemma that we will use in conjunction with Lemma˜12.
@@
$\dim{S_{\mathfrak{p}}}-\dim{(S/\mathfrak{b})_{\mathfrak{p}}}=\dim{R_{\mathfrak{p}}}-\dim{(R/\mathfrak{p}^{\prime})_{\mathfrak{p}}}.$
@@
For the ideal $\mathfrak{a}:=R\cap\mathfrak{b}$, we have a finite extension $R/\mathfrak{a}\hookrightarrow S/\mathfrak{b}$, and hence a finite extension $(R/\mathfrak{a})_{\mathfrak{p}}\hookrightarrow(S/\mathfrak{b})_{\mathfrak{p}}$ for any prime ideal $\mathfrak{p}\subset R$. In particular, $\dim{(S/\mathfrak{b})_{\mathfrak{p}}}=\dim{(R/\mathfrak{a})_{\mathfrak{p}}}$ and $\dim{S_{\mathfrak{p}}}=\dim{R_{\mathfrak{p}}}$. Next, note that $\mathrm{Supp}_{R}(S/\mathfrak{b})=\mathrm{Supp}_{R}(R/\mathfrak{a})=\{\mathfrak{p}\in\operatorname{Spec}(R)\mid\mathfrak{a}\subseteq\mathfrak{p}\}.$ Indeed, the first equality again follows from the observation that $(R/\mathfrak{a})_{\mathfrak{p}}\hookrightarrow(S/\mathfrak{b})_{\mathfrak{p}}$ is an extension, and the second equality is elementary. So the given prime ideal $\mathfrak{p}\in\mathrm{Supp}_{R}(S/\mathfrak{b})$ contains $\mathfrak{a}$, and $\dim{(R/\mathfrak{a})_{\mathfrak{p}}}=\dim{(R/\mathfrak{p}^{\prime})_{\mathfrak{p}}}$ for some prime ideal $\mathfrak{p}^{\prime}$ of $R$ minimal among the primes satisfying $\mathfrak{a}\subseteq\mathfrak{p}^{\prime}\subseteq\mathfrak{p}$. In particular, $\mathfrak{p}^{\prime}\in\mathrm{Supp}_{R}(S/\mathfrak{b})$. ∎
@@ 1.4 Good groups and Steinberg’s Theorem
    The representation ring $\mathrm{R}G$ is a tensor product of a polynomial ring and a Laurent ring, i. e., isomorphic to the ring $\mathbb{Z}[x_{1},\dots,x_{m},y_{1}^{\pm 1},\dots,y_{n}^{\pm 1}]$ for certain $m$ and $n$.
@@
To avoid endlessly repeating these somewhat intricate conditions, we simply define a good Lie group to be a compact connected Lie group $G$ that satisfies the equivalent conditions of Theorem˜14. The result stated as Steinberg’s Theorem is just the implication (a $\Rightarrow$ c) of this theorem combined with the following observation:
@@
Suppose $G$ is compact connected with torsion-free fundamental group. Writing $G$ as $(G^{\prime}\times T)/\Gamma$ as above, we find that the fundamental group of $G^{\prime}$ is also torsion-free. As $G^{\prime}$ is semisimple, this implies that $G^{\prime}$ is simply-connected, and hence a direct product of simply-connected simple groups. So $G$ satisfies condition (a) of Theorem˜14. ∎
Condition (b) in Theorem˜14 implies:
@@
For good $G$, $\mathrm{R}G$ is a localization of a polynomial ring over the integers, which is clearly noetherian and of finite global dimension. ∎
Condition (b) also implies that, for good $G$, the notions of “projective dimension” and “weak/flat dimension” are equivalent for finitely generated $\mathrm{R}G$-modules. Indeed:
@@
The implication (a $\Rightarrow$ b) of Theorem˜14 is fairly elementary, and the implication (b $\Rightarrow$ c), which is the main ingredient in the result stated as Steinberg’s Theorem in the introduction, was in fact proved slightly earlier by Pittie [Pittie:Homogeneous], modulo the above generalization of the Quillen–Suslin theorem. Thus, from today’s vantage point, the main contribution of Steinberg is a proof of the opposite implication: a precise characterization of those compact Lie groups that have the desired property (c). Moreover, Steinberg’s proof includes an explicit basis of $\mathrm{R}H$ as an $\mathrm{R}G$-module.
We can easily sketch Pittie’s short argument for the implication (b $\Rightarrow$ c) based on the results and observations already included above: For a maximal torus $T$ in a good group $G$, both $\mathrm{R}G$ and $\mathrm{R}T$ are regular noetherian rings (Lemma˜16). As $\mathrm{R}T$ is in any case finitely generated as an $\mathrm{R}G$-module (Lemma˜7), it thus follows from [Serre:LA, Chapter III.D, § 5, Theorem 13 (b)] that $\mathrm{R}T$ is a projective $\mathrm{R}G$-module. For a general subgroup of maximal rank $H$, we may assume that $H$ contains $T$, and use holomorphic induction (Theorem˜11) to deduce that $\mathrm{R}H$ is a direct summand of $\mathrm{R}T$. In particular, $\mathrm{R}H$ is also projective as an $\mathrm{R}G$-module. Finally, as we have just seen in Lemma˜17, this is today known to imply that $\mathrm{R}H$ is free.
While Theorem˜14 only deals with subgroups of maximal rank, the following consequence for subgroups of arbitrary ranks was proved in [SCSQ, Corollary 3.10]:
@@
We do not reproduce the short proof of Corollary˜19 here, but we note that besides holomorphic induction (Theorem˜11) the main ingredient is a change-of-rings theorem for Tor. We record here two such theorems that we will have occasion to use later.
@@
This is a special case of [CartanEilenberg, Chapter IX, Theorem 2.8]. Cartan and Eilenberg work with non-commutative $K$-projective algebras $\Lambda$, $\Gamma$, $\Sigma$ over a commutative ring $K$. We have specialized the statement to the case of $\mathbb{Z}$-free commutative $\mathbb{Z}$-algebras. Note that the isomorphism $r$ of [CartanEilenberg, Chapter IX, Proposition 2.1], on which the theorem relies, is an isomorphism of $\Sigma\otimes\Gamma$-algebras in this commutative situation. ∎
@@ 1.5 Supports
In [Segal:RG], Segal analyses the prime spectrum $\operatorname{Spec}(\mathrm{R}G)$ of an arbitrary compact Lie group $G$. In particular, Segal shows that, for each prime ideal $\mathfrak{p}\subset\mathrm{R}G$, there is a subgroup $S_{\mathfrak{p}}$, unique up to conjugation, which is minimal with respect to the property that $\mathfrak{p}$ is the preimage of a prime ideal of $\mathrm{R}(S_{\mathfrak{p}})$ under the restriction $\mathrm{R}G\to\mathrm{R}(S_{\mathfrak{p}})$. This subgroup $S_{\mathfrak{p}}$, or rather any subgroup in this conjugacy class, is referred to as the support of $\mathfrak{p}$. The following proposition explains how Segal’s “supporting subgroups” and the usual notion of “prime ideals supporting a module” are related.
@@
    $\dim{\mathrm{R}G/\mathfrak{p}}\leq\operatorname{rank}H+1.$
@@
Claim (a) is clear from the equivalence (c $\Leftrightarrow$ d) of Proposition˜22.
The first equality in Claim (b) holds in general for finitely generated modules over a commutative ring. See [AtiyahMacDonald, Chapter 3, Exercise 19 (iv)] for the general fact, and Lemma˜7 for finite generation in the case at hand. The second equality in Claim (b) again follows from the equivalence (c $\Leftrightarrow$ d) of Proposition˜22.
For claim (c), recall from [Segal:RG] that the support $S_{\mathfrak{p}}$ is topologically cyclic, i. e., contains an element $s$ whose powers are dense in $S_{\mathfrak{p}}$. This element $s$ is contained in some maximal torus of $G$, so we may as well assume $s\in T$ and hence $S_{\mathfrak{p}}\subseteq T$. The ideal $\mathfrak{p}\subset\mathrm{R}G$ is thus the inverse image of an ideal $\mathfrak{p}_{0}\subset\mathrm{R}(S_{\mathfrak{p}})$ under a restriction map that factors as
@@
Recall that any inclusion of a subgroup into a torus induces a surjection on representation rings (Lemma˜4 (a)). Writing $\mathfrak{p}_{T}\subset\mathrm{R}T$ for the intermediate preimage of $\mathfrak{p}_{0}$, we obtain induced morphisms of quotient rings as follows:
@@
Here, the first map is a finite extension (see Lemma˜9), so $\dim{\mathrm{R}G/\mathfrak{p}}=\dim{\mathrm{R}T/\mathfrak{p}_{T}}$, and we deduce $\dim{\mathrm{R}G/\mathfrak{p}}\leq\dim{\mathrm{R}(S_{\mathfrak{p}})}$. Moreover, by Lemma˜6 (c), $\dim{\mathrm{R}(S_{\mathfrak{p}})}=\operatorname{rank}{S_{\mathfrak{p}}}+1$, and as $S_{\mathfrak{p}}$ is conjugate to a subgroup of $H$ we have $\operatorname{rank}{S_{\mathfrak{p}}}\leq\operatorname{rank}H$.
@@
Let us again write $\mathfrak{p}_{T}$ for the intermediate preimage in $\mathrm{R}T$, which by construction lies over $\mathfrak{p}$. As the Weyl group $\mathrm{W}$ acts trivially on $\mathrm{Z}G$, the restriction $\mathrm{R}T\twoheadrightarrow\mathrm{R}(\mathrm{Z}G)$ is equivariant with respect to the canonical $\mathrm{W}$-action on $\mathrm{R}T$ and the trivial action on $\mathrm{R}(\mathrm{Z}G)$. So $\mathfrak{p}_{T}\subset\mathrm{R}T$ is $\mathrm{W}$-invariant, and it follows from Lemma˜9 (b) that $\mathfrak{p}_{T}$ is the only ideal over $\mathfrak{p}$. ∎
For the biquotient conditions appearing in the following two examples, see the introduction or Definition˜26 immediately below.
Example 24 (Strict biquotient condition).
If $(G,H_{1},H_{2})$ satisfy the strict biquotient condition, then the only prime ideals $\mathfrak{p}$ of $\mathrm{R}G$ supporting both $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ are the prime ideals $\mathfrak{I}G$ and $\mathfrak{I}_{p}G$ from Example˜10. Indeed, Proposition˜23 (b) identifies the intersection $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$ with $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(1)$, and by the equivalence (b $\Leftrightarrow$ c) of Proposition˜22 this latter support consists precisely of the primes containing $\mathfrak{I}G$.
Example 25 (Lax biquotient condition).
If $(G,H_{1},H_{2})$ satisfy the lax biquotient condition, then any prime ideal $\mathfrak{p}$ of $\mathrm{R}G$ supporting both $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ also supports $\mathrm{R}(\mathrm{Z}G)$}. Indeed, by Proposition˜23 (b), any such prime $\mathfrak{p}$ lies in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(H_{1}\cap gH_{2}g^{-1})$ for some $g\in G$, each of the intersections $H_{1}\cap gH_{2}g^{-1}$ is contained in the centre $\mathrm{Z}G$ by assumption, and hence $\mathfrak{p}$ lies in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$ by Proposition˜23 (a).